2011 Ghc1 100 Ghc2 50 2012 Ghc1 200 Ghc2 100 2013 Ghc2 200 Ghc4 100 Base year: 2011 Required: a. Compute for each year the i. nominal GDP, ii. real GDP, (3marks) iii. the GDP deflator. (3marks) b. Compute for 2012 and 2013 from the respective preceding years the percentage change in i. nominal GDP, (ii. real GDP, iii. the GDP deflator (2marks) 2 iv. For each year, identify the variable that does not change. (1mark) Explain in words why your answer makes sense. c. Did economic well-being rise more in 2012 or 2013? Explain.


1
Expert's answer
2021-08-02T11:07:09-0400

1

(a)(I)Nominal GDP=price × quantity

"2011=(1×100)+(2×50)=200\\2012=(1×200)+(2×100)=400\\2013=(2×200)+(4×100)=800"

(ii) real GDP=price of base year × quantity of present year

"2011=(1×100)+(2×50)=200\\2012=(1×200)+(2×100)=400\\2013=(1×200)+(2×100)=400"

(iii)GDP deflator="\frac{nominal\space gdp}{real\space gdp}×100"

"2011=\frac{200}{200}×100=100\\2012=\frac{400}{400}×100=100\\2013=\frac{800}{400}×100=200"

(b) percentage change"=\frac{GDP \space of \space present \space year-gdp \space of \space base\space year}{gdp \space of \space base\space year}×100"

(I) nominal GDP

"2012=\frac{400-200}{200}×100=100\%\\2013=\frac{800-200}{200}×100=300\%"

(ii}real GDP

"2012=\frac{400-200}{200}×100=100\%\\2013=\frac{400-200}{200}×100=100\%"

(iv)GDP deflator

"2012=\frac{100-100}{100}×100=0\%\\2013=\frac{200-100}{100}×100=100\%"

(Iv) quantity does not change for both 2012 and 2013. This is because the calculation is based on the base year.

(c) The economy rose more in 2013 because the nominal GDP percentage rate change is greater in 2013